OpenStudy (anonymous):

Find the midpoint of the following segment created by these pairs of endpoints: (0, 1), (0, 5) 1: (1, 5) 2: (0, 2.5) 3: (0, 3) 4: (5, 1) I am going with c, i think i am starting to under stand it :D

OpenStudy (anonymous):

Find the length of the following segment created by these pairs of endpoints: (1, 2), (7, 2) 7 5 6 2 it would be 2 right?

OpenStudy (anonymous):

the answers number 3

OpenStudy (anonymous):

Yes, you're starting to get it with the midpoints. For the line segment length, it could help (both you and people helping) if you showed a bit more of your reasoning why. Lengths of line segments in coordinate systems are calculated using the Pythagorean theorem.

OpenStudy (anonymous):

first one is (0,3) then the second one is 2

OpenStudy (anonymous):

when you use the numbers to show vs X1 and so on it helped when you say 5 + 6 / 2 i looked at the problem and then i used that for the next problem

OpenStudy (anonymous):

to do this: Find the length of the following segment created by these pairs of endpoints: (4.5, 7), (1.1, 4.3) I would do 4+1 / 3? or (4.5,7) + (1.1,4.3) / 2?

OpenStudy (anonymous):

If you mark the two points in the xy plane and connect them with a straight line segment, you can see that what you're looking for is actually the length of a hypotenuse. The other two sides of the triangle being the difference of coordinates. If the your points were for instance (1,1) and (4,2), then the difference of their x-coordinates is 4-1=4. The difference of their y-coordinates is 2-1=1. Thus you are dealing with a plane triangle with base length 4 and height 1. The hypotenuse of that triangle corresponds to the shortest distance between the given two points. Let's mark the distance between the two points with d, According to the Pythagorean theorem, it now holds that d^2=4^2+1^2. Solve the equation. And do try sketching on grid paper what's happening here. It should really clear things up.

OpenStudy (anonymous):

That is to say: looking for the middle point of a line segment and looking for the length of a line segment are two different tasks with different type of results. Middlepoint is a point, a pair of coordinates in this case. Length is just a bare positive number. You can't solve both problems using the same method. :)

OpenStudy (anonymous):

O jod thats a long answer xD give me a sec to take it in ps i have to go ifor a but

OpenStudy (anonymous):

THank you for the help understanding